Martina008:

given that f(x) =x^2-4x-32 and g(x)=x-8, find f(x)+g(x) and express the result as a polynomial in simpilest form

1st4r:

to find tht you have add the two functions together, f(x) = x^2 - 4x - 32 g (x) = x-8. then you add it tgthr. f(x) + g(x) = (x^2-4x-32) + (x-8) then you needa combine the like terms from f(x) + (gx) = x^2 - 4x + x -32- 8 and u get it answer

Martina008:

thanl you

1st4r:

yeah fs